# Pseudo-Boolean Proof Logging for Optimal Classical Planning

### v1

@@ Conclusion
We believe these certificates to be much more general than ones considered in previous research because they work on a more fundamental level of abstraction, build on an established proof systems that has proved its utility in many other areas of AI, and are able to directly incorporate numerical reasoning. In future work, this generality has to be confirmed by applying these certificates to other optimal planning approaches that current approaches cannot handle. In particular, we believe that pseudo-Boolean constraints can generically handle the concept of cost partitioning (Katz and Domshlak 2010), perhaps the most important optimal planning technique, because they are able to reason numerically at the level of individual state transitions in a way that previously proposed certificates cannot.
@@ Acknowledgments
This research was supported by the Swiss National Science Foundation (SNSF) as part of the project “Unifying the Theory and Algorithms of Factored StateSpace Search” (UTA).
Supplemental Material “Pseudo-Boolean Proof Logging for Optimal Classical Planning”
Proofs for Cost Lemmas
We will use an extended cutting planes proof technique that we did not introduce in the background section, namely redundance-based strengthening (RED): if for a PB constraint $C$ there is a substitution $\rho$ such that $\mathcal{C}\cup\{\lnot C\}\models\mathcal{C}\cup\{\lnot C\}\!\!\upharpoonright_{\rho}$ then we can infer $C$. We only use the simple form with an empty substitution. Since it cannot invalidate the core formula $\mathcal{C}$, we only need to derive that $\mathcal{C}\cup\{\lnot C\}\models\lnot C$ to establish $C$. We can think of this as a more flexible form of a proof by contradiction than simple RUP.
@@
Proof-logging PDB Heuristics: Proofs
@@
Proof-logging $h^{\textup{max}}$: Proofs
@@
Assume (a) $r_{s}\geq 1$, (b) $\textit{cost}_{\geq\max\{0,B-h^{\textup{max}}(s)\}}\geq 1$, and (c) $\overline{r^{\textup{max}}_{s}}\geq 1$. From (a), we get for all $v\in\mathcal{V}\setminus s$ that $\bar{v}\geq 1$, so with (17) we get for all such $v$ that $r_{v}\geq 1$. For all $v\in s$, the heuristic determines $W^{\textup{max}}(v)=V^{\textup{max}}(v)=0$, and thus $B-h^{\textup{max}}(s)+W^{\textup{max}}(v)\leq B$, so from (b) and (17) we get $r_{v}\geq 1$ also for these variables. Together, we derive with (18) that $r^{\textup{max}}_{s}\geq 1$, contradicting (c). ∎
@@
Assume (a) $r_{G}\geq 1$, (b) $\overline{\textit{cost}_{\geq B}}\geq 1$ and (c) $r^{\textup{max}}_{s}\geq 1$. Consider a goal variable $g$ with maximum max value. For this $g$, it holds that $V^{\textup{max}}(g)=h^{\textup{max}}(s)$. From (a) we derive that $g\geq 1$, which gives together with (b) and (17) $\overline{r_{g}}\geq 1$. Thus we get from (18) that $\overline{r^{\textup{max}}_{s}}\geq 1$, contradicting (c). ∎
@@
Then we derive $(r^{\textup{max}}_{s}\land r_{a})\rightarrow r^{\prime}_{v}$ for all variables $v$ by RUP: Assume (h) $r^{\textup{max}}_{s}\geq 1$, (i) $r_{a}\geq 1$, and (j) $\overline{r^{\prime}_{v}}\geq 1$. From (i) and (7) we get (k) $\Delta c^{=\textit{cost}(a)}\geq 1$. From (h) with (17) and (18) we derive (l) $K_{\geq B-h^{\textup{max}}(s)+W^{\textup{max}}(p)}\geq 1$ as in case $W^{\textup{max}}(p)=h^{\textup{max}}(s)$.
If $v\in\textit{del}(a)$, we have from (i) and (7) that $\overline{v^{\prime}}\geq 1$. With (17) this yields $r^{\prime}_{v}\geq 1$, contradicting (j).
If $v\in\textit{add}(a)$, (**) with (l) and (k) to derive $K^{\prime}_{\geq B-h^{\textup{max}}(s)+W^{\textup{max}}(p)+\textit{cost}(a)}$, which together with (****) gives $K^{\prime}_{\geq B-h^{\textup{max}}(s)+W^{\textup{max}}(v)}$. With (17) from $H_{\textup{max}}^{\prime}$, we get $r^{\prime}_{v}\geq 1$, contradicting (j).
If $v\in\mathcal{V}\setminus\textit{evars}(a)$, (j) and (17) yield (m) $v^{\prime}\geq 1$ and (l) $\overline{K^{\prime}_{\geq B-h^{\textup{max}}(s)+W^{\textup{max}}(v)}}\geq 1$. From (i) and (7) we get $eq_{v,v^{\prime}}\geq 1$, which with (m) yields (n) $v\geq 1$. We use (l) with (***) to get $\overline{K^{\prime}_{\geq B-h^{\textup{max}}(s)+W^{\textup{max}}(v)+cost(a)}}\geq 1$, which gives with (**) and (k) that $\overline{K_{\geq B-h^{\textup{max}}(s)+W^{\textup{max}}(v)}}\geq 1$. With (17) and (n) this yields $\overline{r_{v}}\geq 1$, which with (18) gives $\overline{r^{\textup{max}}_{s}}\geq 1$, contradicting (h).
RUP can now derive that $(r^{\textup{max}}_{s}\land r_{a})\rightarrow{r^{\textup{max}}_{s}}^{\prime}$: Assume (o) $r^{\textup{max}}_{s}\geq 1$, (p) $r_{a}\geq 1$ and (q) $\overline{r^{\textup{max}}_{s}}^{\prime}\geq 1$. For all $v\in\mathcal{V}$, we get with (o), (p) and the previously derived $(r^{\textup{max}}_{s}\land r_{a})\rightarrow r^{\prime}_{v}$, that $r^{\prime}_{v}\geq 1$. With (18) from $H_{\textup{max}}$, this yields ${r^{\textup{max}}_{s}}^{\prime}\geq 1$, contradicting (q).
@@
State Set Extension Lemma
@@
Efficiently Proof-Logging Pattern Database Heuristics
@@
Let $\Pi=\langle\mathcal{V},\mathcal{A},I,G\rangle$ be a STRIPS planning task and $P\subseteq\mathcal{V}$. The PB circuit for the PDB heuristic with pattern $P$ is the PB circuit $\langle H_{\textup{PDB}},r_{\textup{PDB}}\rangle$ with input variables $\mathcal{V}\cup\mathcal{V}_{c}$, where $H_{\textup{PDB}}$ contains for each abstract state $s_{\alpha}\subseteq P$ with $d(s_{\alpha})<\infty$ two constraints
@@
$(r_{G}\land\overline{\textit{cost}_{\geq B}})\rightarrow\overline{r_{\textup{PDB}}}$ from $\mathcal{C}_{\textup{goal}}\cup H_{\textup{PDB}}$.
@@
The next lemma again considers individual actions and abstract states, but takes cost into account and allows to derive that the successor situation is covered by the certificate.
@@
The next lemma considers the case that the successor state of some backwards-reachable state is not backwards reachable, and allows to derive that this situation is covered by the certificate.
@@
The following lemma allows us to derive that any successor of any state in a region that is not backwards-reachable, is covered by the certificate:

### v2

@@ Conclusion
We believe these certificates to be much more general than the ones considered in previous research because they work on a more fundamental level of abstraction, build on an established proof systems that has proved its utility in many other areas of AI, and are able to directly incorporate numerical reasoning. In future work, this generality has to be confirmed by applying these certificates to other optimal planning approaches that current approaches cannot handle. In particular, we believe that pseudo-Boolean constraints can generically handle the concept of cost partitioning (Katz and Domshlak 2010), perhaps the most important optimal planning technique, because they are able to reason numerically at the level of individual state transitions in a way that previously proposed certificates cannot.
@@ Acknowledgments
This research was supported by the Swiss National Science Foundation (SNSF) as part of the project “Unifying the Theory and Algorithms of Factored StateSpace Search” (UTA). Jakob Nordström was supported by the Independent Research Fund Denmark grant 9040-00389B.
Appendix A Appendix
The following sections contain additional details not provided in the ICAPS 2025 version of this paper.
Appendix B Proofs for Cost Lemmas
We will use an extended cutting planes proof technique that we did not introduce in the background section, namely redundance-based strengthening (RED): if for a PB constraint $C$ there is a substitution $\rho$ such that $\mathcal{C}\cup\{\lnot C\}\models\mathcal{C}\cup\{C\}\!\!\upharpoonright_{\rho}$ then we can infer $C$. We only use the simple form with an empty substitution. Since it cannot invalidate the core formula $\mathcal{C}$, we only need to derive that $\mathcal{C}\cup\{\lnot C\}\models C$ to establish $C$. We can think of this as a more flexible form of a proof by contradiction than simple RUP.
@@ Appendix B Proofs for Cost Lemmas
Appendix C Proof-Logging PDB Heuristics: Proofs
@@
Appendix D Proof-Logging $h^{\textup{max}}$: Proofs
@@
Assume (a) $r_{s}\geq 1$, (b) $\textit{cost}_{\geq\max\{0,B-h^{\textup{max}}(s)\}}\geq 1$, and (c) $\overline{r^{\textup{max}}_{s}}\geq 1$. From (a), we get for all $v\in\mathcal{V}\setminus s$ that $\bar{v}\geq 1$, so with (17) we get for all such $v$ that $r_{v,s}\geq 1$ (*). For all $v\in s$, the heuristic determines $W^{\textup{max}}(v)=V^{\textup{max}}(v)=0$, and thus $B-h^{\textup{max}}(s)+W^{\textup{max}}(v)\leq B$, so from (b) and (17) we get $r_{v,s}\geq 1$ also for these variables (**).
Assumption (b) corresponds to $K_{\geq B-h^{\textup{max}}(s)}$.
Together with (*) and (**), we derive with (18) that $r^{\textup{max}}_{s}\geq 1$, contradicting (c). ∎
@@
Assume (a) $r_{G}\geq 1$, (b) $\overline{\textit{cost}_{\geq B}}\geq 1$ and (c) $r^{\textup{max}}_{s}\geq 1$.
If $G\neq\emptyset$, consider a goal variable $g$ with maximum max value. For this $g$, it holds that $V^{\textup{max}}(g)=h^{\textup{max}}(s)$. From (a) we derive that $g\geq 1$, which gives together with (b) and (17) $\overline{r_{g}}\geq 1$. Thus we get from (18) that $\overline{r^{\textup{max}}_{s}}\geq 1$, contradicting (c).
If $G=\emptyset$ then $h^{\textup{max}}(s)=0$ for all states $s$ and $W^{\textup{max}}(v)=0$ for all variables $v$. From (c) we get that $K_{\geq B}\geq 1$, contradicting (b). ∎
@@
The overall argument of the proof is as follows:
If the action has a precondition $p$, we can conclude that $p\geq 1$, which allows us to derive with (17) that $K_{\geq B-h^{\textup{max}}(s)+W^{\textup{max}}(p)}\geq 1$. We can establish $K^{\prime}_{\geq B-h^{\textup{max}}(s)}$ using lemmas Proof-Logging Heuristic Search and Proof-Logging Heuristic Search.
To establish $r^{\prime}_{v,s}\geq 1$ for all variables, we consider the precondition with maximal $W^{\textup{max}}$ value.
If $W^{\textup{max}}(p)=h^{\textup{max}}(s)$, then $K_{\geq B-h^{\textup{max}}(s)+W^{\textup{max}}(p)}\geq 1$ corresponds to $K_{\geq B}\geq 1$. We derive $K^{\prime}_{\geq B}\geq 1$, which is sufficient to establish $r^{\prime}_{v,s}\geq 1$.
If $W^{\textup{max}}(p)<h^{\textup{max}}(s)$, we establish $r^{\prime}_{v,s}\geq 1$ depending on whether and how $v$ occurs in the effect of $a$. If it is a delete effect, then $\overline{v^{\prime}}\geq 1$. If it is an add effect, we exploit that $p$ has maximal $W^{\textup{max}}$ value among the preconditions, so from the definition of the maximum heuristic we can see that $W^{\textup{max}}(v)\leq W^{\textup{max}}(p)+\textit{cost}(a)$. Establishing $K^{\prime}_{\geq B-h^{\textup{max}}(s)+W^{\textup{max}}(p)+\textit{cost}(a)}$ and using this insight gives $K^{\prime}_{\geq B-h^{\textup{max}}(s)+W^{\textup{max}}(v)}\geq 1$. If $v\notin\textit{evars}(a)$, we can carry over the reason for $r_{v,s}\geq 1$ to the primed variables.
If the precondition of the action is empty, we build on $K_{\geq B-h^{\textup{max}}(s)}$, which gives $K^{\prime}_{\geq B-h^{\textup{max}}(s)+\textit{cost}(a)}$ and $K^{\prime}_{\geq B-h^{\textup{max}}(s)}$ using lemmas Proof-Logging Heuristic Search and Proof-Logging Heuristic Search. For establishing $r^{\prime}_{v,s}\geq 1$ for all variables, we proceed as before for case $W^{\textup{max}}(p)<h^{\textup{max}}(s)$ (replacing $W^{\textup{max}}(p)$ with 00). For the special case $h^{\textup{max}}(s)=0$, we can instead use the shorter argument as in the previous case $W^{\textup{max}}(p)=h^{\textup{max}}(s)$, building on $K_{\geq B}\geq 1$.
We not turn to specifying the details of this overall argument:
@@
Then we derive $(r^{\textup{max}}_{s}\land r_{a})\rightarrow r^{\prime}_{v,s}$ for all variables $v$ by RUP: Assume (h) $r^{\textup{max}}_{s}\geq 1$, (i) $r_{a}\geq 1$, and (j) $\overline{r^{\prime}_{v,s}}\geq 1$. From (i) and (7) we get (k) $\Delta c^{=\textit{cost}(a)}\geq 1$. From (h) with (17) and (18) we derive (l) $K_{\geq B-h^{\textup{max}}(s)+W^{\textup{max}}(p)}\geq 1$ as in case $W^{\textup{max}}(p)=h^{\textup{max}}(s)$.
If $v\in\textit{del}(a)$, we have from (i) and (7) that $\overline{v^{\prime}}\geq 1$. With (17) this yields $r^{\prime}_{v,s}\geq 1$, contradicting (j).
If $v\in\textit{add}(a)$, we use (**) with (l) and (k) to derive $K^{\prime}_{\geq B-h^{\textup{max}}(s)+W^{\textup{max}}(p)+\textit{cost}(a)}$, which together with (****) gives $K^{\prime}_{\geq B-h^{\textup{max}}(s)+W^{\textup{max}}(v)}$. With (17) from $H_{\textup{max},s}^{\prime}$, we get $r^{\prime}_{v,s}\geq 1$, contradicting (j).
If $v\in\mathcal{V}\setminus\textit{evars}(a)$, (j) and (17) yield (m) $v^{\prime}\geq 1$ and (n) $\overline{K^{\prime}_{\geq B-h^{\textup{max}}(s)+W^{\textup{max}}(v)}}\geq 1$. From (i) and (7) we get $eq_{v,v^{\prime}}\geq 1$, which with (m) yields (o) $v\geq 1$. We use (n) with (***) to get $\overline{K^{\prime}_{\geq B-h^{\textup{max}}(s)+W^{\textup{max}}(v)+cost(a)}}\geq 1$, which gives with (**) and (k) that $\overline{K_{\geq B-h^{\textup{max}}(s)+W^{\textup{max}}(v)}}\geq 1$. With (17) and (o) this yields $\overline{r_{v,s}}\geq 1$, which with (18) gives $\overline{r^{\textup{max}}_{s}}\geq 1$, contradicting (h).
[…]
RUP can now derive that $(r^{\textup{max}}_{s}\land r_{a})\rightarrow{r^{\textup{max}}_{s}}^{\prime}$: Assume (v) $r^{\textup{max}}_{s}\geq 1$, (w) $r_{a}\geq 1$ and (x) $\overline{{r^{\textup{max}}_{s}}^{\prime}}\geq 1$. For all $v\in\mathcal{V}$, we get with (v), (x) and the previously derived $(r^{\textup{max}}_{s}\land r_{a})\rightarrow r^{\prime}_{v,s}$, that $r^{\prime}_{v,s}\geq 1$. We use (v), (x) and the previously derived $(r^{\textup{max}}_{s}\land r_{a})\rightarrow K^{\prime}_{\geq B-h^{\textup{max}}(s)}$ to get $K^{\prime}_{\geq B-h^{\textup{max}}(s)}\geq 1$. With (18) from $H_{\textup{max},s}^{\prime}$, these yield ${r^{\textup{max}}_{s}}^{\prime}\geq 1$, contradicting (x).
@@
Appendix E State Set Extension Lemma
@@
Appendix F Efficiently Proof-Logging Pattern Database Heuristics
@@
Let $\Pi=\langle\mathcal{V},\mathcal{A},I,G\rangle$ be a STRIPS planning task and $P\subseteq\mathcal{V}$. The PB circuit for the PDB heuristic with pattern $P$ is the PB circuit $\langle H_{\textup{PDB}},r_{\textup{PDB}}\rangle$ with input variables $\mathcal{V}\cup\mathcal{V}_{c}$, where the set represented by $H_{\textup{PDB}}$ contains for each abstract state $s_{\alpha}\subseteq P$ with $d(s_{\alpha})<\infty$ two constraints
@@
$(r_{G}\land{r_{\textup{PDB}}})\rightarrow{\textit{cost}_{\geq B}}$ from $\mathcal{C}_{\textup{goal}}\cup H_{\textup{PDB}}$.
@@
The next lemma again considers individual actions and abstract states, but takes cost into account and allows to derive that the successor state-cost pair is contained in the invariant of the certificate.
@@
The next lemma considers the case that the successor state of some backwards-reachable state is not backwards reachable, and allows to derive that this state-cost pair is contained in the invariant of the certificate.
@@
The following lemma allows us to derive that any successor of any state in a region that is not backwards-reachable, is contained in the invariant: